OpenStudy (anonymous):

Calculate the circumference of the circles with the following dimensions. C = circumference D = diameter R = radius PI = 3.14 R = 9 ft R = 11 ft R = 7 ft

OpenStudy (anonymous):

Cirumference C=2*PI*R or C=PI*D (Diameter =2*Radius)

OpenStudy (anonymous):

use this to find out circumference C=2*3.14*9=56.52 ft Similarly you can find for other R where R=11ft, R=7ft

OpenStudy (anonymous):

The equation to find the circumference is: \[C=2 \pi R\] Therefore: R(7): \[2*\pi * 7 = 14\pi\] R(9): \[2*\pi * 9 = 18\pi\] R(11): \[2*\pi * 11 = 21\pi\]